The width of a rectangular room is 3 feet shorter than it is long. If the perimeter of the room is 42 feet, what is the width of the room?
1. 12 feer
2. 9 feet
3. 18 feet
4. 6 feet

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

Same problem, different numbers.
==========================
A rectangular room is 7 feet shorter than it is wide. If the perimeter of the room is 70 feet, what is the length of the room?
------------------
L - W = 7
2L + 2W = 70

Answer by ikleyn(52781)

.
The correct system of equations is


    L - W =  3      (1)

    L + W = 21      (2)     ( <<<---===  21 = 42/2 (!) )



Add equations (1) and (2), and you will get  2L = 3 + 21 = 24  ====>  L = 24/2 = 12.


Then from equation (2),  W = 21 - 12 = 9.



ANSWER.  The width of the room is 9 feet.